OpenStudy (anonymous):

A segment 12" long is divided into two segments having lengths in the ratio of 2:3. Find the length of the shortest segment. 2 2 2/5 4 4/5 7 1/5

Parth (parthkohli):

You set up an equation here. Easy. \(\Large \color{MidnightBlue}{\Rightarrow 2x + 3x = 12 }\) \(\Large \color{MidnightBlue}{\Rightarrow 5x = 12 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = {12 \over 5} = 2{2 \over 5} }\)

Parth (parthkohli):

shorter segment is 2 * 12/5 = 24/5

Parth (parthkohli):

Same as 4 4/5

OpenStudy (anonymous):

in such cases add the ratio first : viz. 2+3 = 5 then use this step : first part : (2/5)*12 next : (3/5)*12 You will get ur answer.

Parth (parthkohli):

@sachinrajsharma it's the shorter segment, actually.

OpenStudy (anonymous):

so then 4 4/5?

OpenStudy (anonymous):

u can then compare the numerator as when u have the same denominator in this case denominator is 5 both cases